Trico Company set the following standard unit costs for its single product.

Direct materials (30 Ibs. @ $4.00 per Ib.) $ 120.00

Direct labor (7 hrs. @ $14 per hr.) 98.00

Factory overhead—Variable (7 hrs. @ $7 per hr.) 49.00

Factory overhead—Fixed (7 hrs. @ $11 per hr.) 77.00

Total standard cost $ 344.00

The predetermined overhead rate is based on a planned operating volume of 80% of the productive capacity of 63,000 units per quarter. The following flexible budget information is available.

Operating Levels 70% 80% 90% Production in units 44,100 50,400 56,700 Standard direct labor hours 308,700 352,800 396,900

Budgeted overhead Fixed factory overhead $ 3,880,800 $ 3,880,800 $ 3,880,800 Variable factory overhead $ 2,160,900 $ 2,469,600 $ 2,778,300 During the current quarter, the company operated at 90% of capacity and produced 56,700 units of product; actual direct labor totaled 393,900 hours. Units produced were assigned the following standard costs.

Direct materials (1,701,000 Ibs. @ $4.00 per Ib.) $ 6,804,000

Direct labor (396,900 hrs. @ $14 per hr.) 5,556,600

Factory overhead (396,900 hrs. @ $18 per hr.) 7,144,200 Total standard cost $ 19,504,800

Actual costs incurred during the current quarter follow. Direct materials (1,686,000 Ibs. @ $5.10 per lb.) $ 8,598,600 Direct labor (393,900 hrs. @ $12.00 per hr.) 4,726,800 Fixed factory overhead costs 3,380,000 Variable factory overhead costs 3,164,200 Total actual costs $ 19,869,600 Problem 08-4A Computing materials, labor, and overhead variances LO P3, P4

Required: 1. Compute the direct materials cost variance, including its price and quantity variances.

2. Compute the direct labor cost variance, including its rate and efficiency variances.

3. Compute the overhead controllable and volume variances.
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1686000 $5.10 $8,598,600.00 1686000 $4.00 $6,744,000.00 1701000 $4.00 $6,804,000.00
$1,854,600.00 U $60,000.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $1,854,600.00 U
Direct material quantity variance $60,000.00 F
Direct material cost variance $1,794,600.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
393900 $12.00 $4,726,800.00 393900 $14.00 $5,514,600.00 396900 $14.00 $5,556,600.00
$787,800.00 F $42,000.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $787,800.00 F
Direct Labor Efficiency variance $42,000.00 F
Direct labor cost variance $829,800.00 F

Solution 3:

Controllable Variance
Actual overhead $6,544,200.00
Budgeted overhead $6,659,100.00
Controllable variance $114,900.00 F
Fixed overhead volume variance
Budgeted fixed overhead $3,880,800.00
Fixed overhead cost applied $4,365,900.00
Fixed overhead volume variance $485,100.00 F